Two airplanes leave an airport at the same
time. The velocity of the first airplane is
710 m/h at a heading of 58◦
. The velocity of
the second is 630 m/h at a heading of 181◦
.
How far apart are they after 2.9 h?
Answer in units of m.
1
Expert's answer
2019-10-04T13:01:42-0400
"d_1=(710)(2.9)=2059\ m"

"d_2=(630)(2.9)=1827\ m"

"d=\sqrt{2059^2+1827^2-2(2059)(1827)\cos{(181-58)\degree}}"

"d=3417\ m"